2019, 200 KGs of oranges are produced at Rs.40/kg, 500 KGs of apples are produced at 120/kg, and 200 Glasses of orange juice are produced at Rs.150/glass.

In 2020, 220 KGs of oranges are produced at Rs.50/kg, 450 KGs of apples are produced at 140/kg, and 230 glasses of orange juice are produced at Rs.135/glass.

Assuming 2019 is the base year; calculate the nominal GDP in 2019



1
Expert's answer
2021-05-05T13:36:26-0400

Solution:

Nominal GDP = Sum of (this year prices x this year quantities)

 

Nominal GDP 2019:

Oranges = 40 "\times" 200 = 8,000

Apples = 120 "\times" 500 = 60,000

Orange juice = 200 "\times" 150 = 30,000

Total = 8,000 + 60,000 + 30,000 = 98,000

Nominal GDP 2019 = 98,000


N/B:

Since 2019 is the base year, Nominal GDP will be the same as Real GDP.

Real GDP 2019 = 98,000